# HAIC: Improving Human Action Understanding and Generation with Better Captions for Multi-modal Large Language Models

### v1

@@ Abstract
Recent Multi-modal Large Language Models (MLLMs) have made great progress in video understanding. However, their performance on videos involving human actions is still limited by the lack of high-quality data. To address this, we introduce a two-stage data annotation pipeline. First, we design strategies to accumulate videos featuring clear human actions from the Internet. Second, videos are annotated in a standardized caption format that uses human attributes to distinguish individuals and chronologically details their actions and interactions. Through this pipeline, we curate two datasets, namely HAICTrain and HAICBench. HAICTrain comprises 126K video-caption pairs generated by Gemini-Pro and verified for training purposes. Meanwhile, HAICBench includes 500 manually annotated video-caption pairs and 1,400 QA pairs, for a comprehensive evaluation of human action understanding. Experimental results demonstrate that training with HAICTrain not only significantly enhances human understanding abilities across 4 benchmarks, but can also improve text-to-video generation results. Both the HAICTrain and HAICBench are released at https://huggingface.co/datasets/KuaishouHAIC/HAIC.
@@
Xiao Wang $1$Equal contribution. [^2] Work done as an intern at Kuaishou Technology.  Jingyun Hua $1$Equal contribution.  Weihong Lin $1$Equal contribution.  Yuanxing Zhang 1 Fuzheng Zhang 1Jianlong WuDi Zhang 1Liqiang Nie  [^3] Corresponding author. 1Kuaishou Technology
@@ 1 Introduction
Multi-modal large language models have notably showcased their preeminence across various video understanding tasks Chen et al. (2024d); Li et al. (2024a); Wang et al. (2024a); Zhang et al. (2024b). Among these, human action understanding plays a critical role in many downstream applications, e.g., human-computer interaction Hayes (2011), autonomous driving Xu et al. (2021), embodied intelligence Gupta et al. (2021), and human video generation Wang et al. (2024b).
A recent study, ShareGPT4Video Chen et al. (2024a) has demonstrated that high-quality and detailed video captions can improve MLLMs’ performance in video understanding. However, most existing works Soomro et al. (2012); Xu et al. (2016); Krishna et al. (2017); Wang et al. (2019); Chen et al. (2024c) provide only coarse captions for human actions, insufficient for understanding fine-grained behaviors. MotionLLM Chen et al. (2024b) introduces the MoVid dataset with fine-grained action captions from MotionX Lin et al. (2023). Nevertheless, this dataset mainly focused on single-person scenarios. For multi-person situations, MoVid only considered consistent group activities like “a group of people performing the Korean dance”. A comprehensive dataset is essential to enhance MLLMs’ understanding of detailed human actions and interactions in both single- and multi-person contexts, critical for tasks like emotional analysis, motivation prediction, and relationship modeling.
@@
To address the above challenges, we propose a novel data generation pipeline composed of two stages. In the video accumulation stage, we accumulate videos from various domains featuring clear, meaningful human actions and identify their specific timestamps. This process is highly selective, with only about 1% of videos meeting our quality criteria after applying various strategies. In the caption annotation stage, we define a caption format that uses human attributes to distinguish individuals and chronologically annotates detailed body actions and interactions for each person (see Figure 1). With this pipeline, we curate two datasets: HAICTrain (Human Action and Interaction Comprehension Training set) and HAICBench (a benchmark for evaluation). HAICTrain contains 126K videos accumulated from WebVid Bain et al. (2021), annotated in our defined caption format by Gemini-1.5-Pro Team et al. (2024). HAICBench includes 500 YouTube videos with human-annotated captions in the same format. Furthermore, we generate 1,400 multiple-choice QA pairs across five categories—human interaction, action detail, action sequence, count, and human attribute—by prompting GPT-4o Hurst et al. (2024) and Gemini-1.5-Pro. Note that all the above machine annotation results undergo review and refinement by human annotators. Experimental results indicate that utilizing HAICTrain for training can remarkably enhance the model’s human action understanding ability by 1.4%-2.1%. Additionally, in MovieGenBench Polyak et al. (2024), our post-trained model surpasses the original model in GSB score of 2.15 and 6.81 in HunyuanVideo Kong et al. and Wanx2.1 Aliyun (2023), respectively.
@@
  • We introduce two datasets: HAICTrain which includes 126K generated-then-verified high-quality video-caption pairs for training; and HAICBench comprising 500 human annotated video-caption pairs and 1,400 QA pairs, designed to evaluate MLLMs’ human action understanding comprehensively.
@@ 2.1 Video Caption Datasets
Most existing video captioning datasets prioritized general video understanding Xu et al. (2016); Zhou et al. (2018); Wang et al. (2019); Miech et al. (2019); Bain et al. (2021); Zellers et al. (2021); Lei et al. (2021); Xue et al. (2022); Chen et al. (2024a, c); Wang et al. (2024d, c); Xiong et al. (2024); Wang et al. (2024b). These datasets included only a subset of human videos, and the action captions tend to be coarse. Some datasets specifically focused on human-centric captioning, including ActivityNet-Captions Krishna et al. (2017), Ego4D Grauman et al. (2022), Kinetic-GEB Wang et al. (2022), ActionHub Zhou et al. (2024) and OpenHumanVid Li et al. (2024b).
@@ 2.2 Video Understanding Benchmarks
Traditional video understanding benchmarks have primarily honed in a fixed set of classes Soomro et al. (2012); Shahroudy et al. (2016); Carreira and Zisserman (2017); Goyal et al. (2017). Recently, there has been a shift towards open-set video understanding, including video captioning and question answering Xu et al. (2016); Wang et al. (2019); Xiao et al. (2021); Li et al. (2024c). While benchmarks like NeXT-QA Xiao et al. (2021), MVBench Li et al. (2024c), and PerceptionTest Patraucean et al. (2023) include action-related questions, they primarily focus on general action recognition rather than fine-grained details. TemporalBench Cai et al. (2024) evaluates event-level action understanding using a different protocol that requires distinguishing correct captions from negative ones. Action-specific benchmarks, such as ActivityNet-Captions Krishna et al. (2017), TGIF-QA Jang et al. (2019), MoVid-Bench Chen et al. (2024b), EGOBody Zhang et al. (2022), and GRAB Taheri et al. (2020), address temporal sequences or human-object interactions but often overlook human-human interactions.
Our HAICBench provides a more comprehensive evaluation of human action understanding, covering five key aspects detailed in Table 1.
@@ 2.3 Multi-modal Large Language Models
MLLMs generally process multi-modal input information and generate language output. Kosmos Huang et al. (2023) introduced an end-to-end framework that integrated visual inputs with LLM from cohesive training. Flamingo Alayrac et al. (2022) and InstructBLIP Dai et al. (2023) merged visual and linguistic features through cross-attention and a Q-Former module, respectively. MiniGPT-4 Zhu et al. (2024) and LLaVA Liu et al. (2023) simplified the integration by linearly projecting visual features directly into the LLM embedding space.
@@ 3.2 HAICTrain
We chose the WebVid-10M dataset Bain et al. (2021) as the video source for training due to its large scale and high vision quality. Initially, we apply the action video accumulation pipeline in Section 3.1 to collect action videos from the WebVid-10M. This process results in a collection of 126K videos, representing 1.2% of the original dataset. Then, we employ Gemini-1.5-Pro Team et al. (2024) to generate captions in the standardized format as depicted in Figure 1 referring to the videos and original captions. The specific prompts used for this process are detailed in Appendix A. We then employ an additional judgement to filter out failure cases that do not follow the pre-defined format or display low quality. This judgement ensures the quality of our data in HAICTrain.
@@ 3.3 HAICBench
Video Caption Annotation. To avoid potential overlap with public benchmarks, we choose the newly proposed Omega-multimodal dataset OMEGA Labs (2024) as our benchmark video source, which comprises over 30 million 2-minute video clips. We apply our video accumulation stage in Section 3.1 to obtain some video clips, and manually check these clips to obtain 500 video clips. Then, human annotators are required to describe the number of human subjects, noting static attributes (gender, age, clothing, accessories) for subject identification, followed by body movements, action sequences, and interactions with others in chronological order.
@@
QA Pair Generation. Based on captions above, we prompt GPT-4o Hurst et al. (2024) and Gemini-1.5-Pro Team et al. (2024) to generate multiple-choice question-answer pairs about human interactions, action details, action sequences, count, and human attributes. The prompts share a similar format with those in MVBench Li et al. (2024c). See Appendix C for details. Each question-answer pair is checked by 2 annotators and will be corrected if there are any mistakes. All options are shuffled to avoid potential bias.
We get a total of 500 video-caption pairs and 1,400 QA pairs for our HAICBench dataset. Detailed statistics are presented in Figure 3. Although each video clip focuses on a single scene and is relatively short (less than 20 seconds), the captions are highly detailed, often exceeding 100 words. The word cloud analysis reveals that our captions provide comprehensive descriptions of actions.
@@ 4 Experiments
Gemini-1.5-Pro Team et al. (2024) - 41.5 29.1 28.7 36.3 54.7 58.7
GPT-4o Hurst et al. (2024) 50* 40.3 32.6 28.3 38.0 52.0 50.7
VideoLLaMA2-7B Cheng et al. (2024) 64 21.5 11.7 18.0 14.7 39.3 24.0
LongVA-7B Zhang et al. (2024a) 64 16.2 11.4 13.0 11.0 25.3 20.3
InternVL2-8B Chen et al. (2024d) 64 29.1 20.0 18.3 22.7 43.3 41.0
Qwen2VL-7B Wang et al. (2024a) 64 28.7 16.6 21.7 21.3 44.7 39.3
LLaVA-Video-7B Zhang et al. (2024b) 64 30.7 19.1 16.0 25.3 44.0 49.0
LLaVA-Video-ActionPro-7B 64 35.7 22.3 17.0 31.7 53.3 54.3
@@
Gemini-1.5-Pro Team et al. (2024) - 66.9 60.3 49.3 75.7 66.7 82.7
GPT-4o Hurst et al. (2024) 50* 60.6 59.1 43.4 64.0 54.0 82.3
VideoLLaMA2-7B Cheng et al. (2024) 64 49.7 35.4 37.7 48.3 62.7 64.3
LongVA-7B Zhang et al. (2024a) 64 50.3 43.1 39.3 56.0 37.3 75.7
InternVL2-8B Chen et al. (2024d) 64 58.2 41.4 40.0 61.7 68.0 80.0
Qwen2VL-7B Wang et al. (2024a) 64 64.4 61.1 44.7 65.3 71.3 79.7
LLaVA-Video-7B Zhang et al. (2024b) 64 64.3 53.7 43.7 67.3 70.7 86.3
LLaVA-Video-ActionPro-7B 64 66.4 56.6 45.0 68.7 76.0 86.0
@@
LLaVA-Video 59.4 58.1 61.8 64.3
60.0 58.3 63.8 64.7
62.1 59.9 65.2 66.4
58.8 55.8 62.1 62.2
LLaVA-OneVision 57.8 47.9 57.1 61.6
59.3 55.7 61.8 62.2
60.6 56.9 62.7 64.1
58.8 55.8 60.4 61.8
@@ 4.1 Experimental Settings
Evaluation dataset. To evaluate human action understanding ability, we executed experiments on several action-related benchmarks, including MVBench, ActivityNet-QA, PerceptionTest, and our HAICBench. Note that we focused on questions related to human actions in these benchmarks. For MVBench, we performed a comparison of all sub-tasks whose names contain “Action”, resulting in 7 types: Action Antonym, Action Count, Action Sequence, Action Prediction, Action Localization, Fine-grained Action, and Unexpected Action. For PerceptionTest, we also selected all questions whose tags are related to action. More details are presented in Appendix B.
Evaluation metrics. For multiple-choice questions in MVBench, PerceptionTest, and HAICBench, we followed the prompting approach in LLaVA-1.5 Liu et al. (2024) and used accuracy as the metric. For ActivityNet-QA which features open-ended questions, we followed the evaluation protocol in Video-ChatGPT Maaz et al. (2024), utilizing GPT-4o-0513 to calculate accuracy. In terms of HAICBench, we implemented two evaluation settings: (1) Standard Evaluation, where the video and question are directly input into an MLLM to generate an answer; and (2) Caption Evaluation, where the model generates a caption for each video, and an LLM (Gemini-1.5-Pro) answers questions based on this caption, thereby assessing caption ability based on QA accuracy. To avoid the LLM guessing answers when captions do not mention related contents, we asked Gemini-1.5-Pro to refuse to answer those questions (see Appendix D).
Implementation Details. We used LLaVA-Video-7B as our initial model and fine-tuned it on the training dataset. All parameters were updated during training. The model was fine-tuned for one epoch with a learning rate of 1e-5 for the LLM and 2e-6 for the vision encoder, with 256 batch size. We sampled 64 frames uniformly in both the training and evaluation stages. During generation, we take the greedy search without randomness. Our experiments utilized 128 NVIDIA A800-80GB GPUs.

### v2

@@ Abstract
Recent Multi-modal Large Language Models (MLLMs) have made great progress in video understanding. However, their performance on videos involving human actions is still limited by the lack of high-quality data. To address this, we introduce a two-stage data annotation pipeline. First, we design strategies to accumulate videos featuring clear human actions from the Internet. Second, videos are annotated in a standardized caption format that uses human attributes to distinguish individuals and chronologically details their actions and interactions. Through this pipeline, we curate two datasets, namely HAICTrain and HAICBench. HAICTrain comprises 126K video-caption pairs generated by Gemini-Pro and verified for training purposes. Meanwhile, HAICBench includes 412 manually annotated video-caption pairs and 2,000 QA pairs, for a comprehensive evaluation of human action understanding. Experimental results demonstrate that training with HAICTrain not only significantly enhances human understanding abilities across 4 benchmarks, but can also improve text-to-video generation results. Both the HAICTrain and HAICBench are released at https://huggingface.co/datasets/KuaishouHAIC/HAIC.
@@
Xiao Wang1111Equal contribution. [^2] Work done as an intern at Kuaishou Technology.  Jingyun Hua1111Equal contribution.  Weihong Lin1111Equal contribution.  Yuanxing Zhang1 Fuzheng Zhang1Jianlong Wu2333Corresponding author.Di Zhang1Liqiang Nie2333Corresponding author. 1Kuaishou Technology  2Shandong University scz.wangxiao@gmail.com, huajingyun@hotmail.com, lwher1996@outlook.com, longo@pku.edu.cn zhfzhkris@outlook.com, jlwu1992@pku.edu.cn, zhangdi08@kuaishou.com, nieliqiang@gmail.com
@@ 1 Introduction
Multi-modal large language models have notably showcased their preeminence across various video understanding tasks Chen et al. (2024e); Li et al. (2024a); Wang et al. (2024a); Zhang et al. (2024b). Among these, human action understanding plays a critical role in many downstream applications, e.g., human-computer interaction Hayes (2011), autonomous driving Xu et al. (2021), embodied intelligence Gupta et al. (2021), and human video generation Wang et al. (2024b).
A recent study, ShareGPT4Video Chen et al. (2024a) has demonstrated that high-quality and detailed video captions can improve MLLMs’ performance in video understanding. However, most existing works Soomro et al. (2012); Xu et al. (2016); Krishna et al. (2017); Wang et al. (2019); Chen et al. (2024d) provide only coarse captions for human actions, insufficient for understanding fine-grained behaviors. MotionLLM Chen et al. (2024b) introduces the MoVid dataset with fine-grained action captions from MotionX Lin et al. (2023). Nevertheless, this dataset mainly focused on single-person scenarios. For multi-person situations, MoVid only considered consistent group activities like “a group of people performing the Korean dance”. A comprehensive dataset is essential to enhance MLLMs’ understanding of detailed human actions and interactions in both single- and multi-person contexts, critical for tasks like emotional analysis, motivation prediction, and relationship modeling.
@@
To address the above challenges, we propose a novel data generation pipeline composed of two stages. In the video accumulation stage, we accumulate videos from various domains featuring clear, meaningful human actions and identify their specific timestamps. This process is highly selective, with only about 1% of videos meeting our quality criteria after applying various strategies. In the caption annotation stage, we define a caption format that uses human attributes to distinguish individuals and chronologically annotates detailed body actions and interactions for each person (see Figure 1). With this pipeline, we curate two datasets: HAICTrain (Human Action and Interaction Comprehension Training set) and HAICBench (a benchmark for evaluation). HAICTrain contains 126K videos accumulated from WebVid Bain et al. (2021), annotated in our defined caption format by Gemini-1.5-Pro Team et al. (2024). HAICBench includes 412 YouTube videos with human-annotated captions in the same format. Furthermore, we generate 2,000 multiple-choice QA pairs across five categories—human interaction, action detail, action sequence, count, and human attribute—by prompting GPT-4o Hurst et al. (2024) and Gemini-1.5-Pro. Note that all the above machine annotation results undergo review and refinement by human annotators. Experimental results indicate that utilizing HAICTrain for training can remarkably enhance the model’s human action understanding ability by 1%-2%. Additionally, in MovieGenBench Polyak et al. (2024), our post-trained model surpasses the original model in GSB score of 2.15 and 6.81 in HunyuanVideo Kong et al. and Wanx2.1 Aliyun (2025), respectively.
@@
  • We introduce two datasets: HAICTrain which includes 126K generated-then-verified high-quality video-caption pairs for training; and HAICBench comprising 412 human annotated video-caption pairs and 2,000 QA pairs, designed to evaluate MLLMs’ human action understanding comprehensively.
@@ 2.1 Video Caption Datasets
Most existing video captioning datasets prioritized general video understanding Xu et al. (2016); Zhou et al. (2018); Wang et al. (2019); Miech et al. (2019); Bain et al. (2021); Zellers et al. (2021); Lei et al. (2021); Xue et al. (2022); Chen et al. (2024a, d); Wang et al. (2024d, 2025b); Xiong et al. (2024); Wang et al. (2024b). These datasets included only a subset of human videos, and the action captions tend to be coarse. Some datasets specifically focused on human-centric captioning, including ActivityNet-Captions Krishna et al. (2017), Ego4D Grauman et al. (2022), Kinetic-GEB Wang et al. (2022), ActionHub Zhou et al. (2024) and OpenHumanVid Li et al. (2024b).
@@ 2.2 Video Understanding Benchmarks
Traditional video understanding benchmarks have primarily honed in on a fixed set of classes Soomro et al. (2012); Shahroudy et al. (2016); Carreira and Zisserman (2017); Goyal et al. (2017). Recently, there has been a shift towards open-set video understanding, including video captioning and question answering Xu et al. (2016); Wang et al. (2019); Xiao et al. (2021); Li et al. (2024c). While benchmarks like NeXT-QA Xiao et al. (2021), MVBench Li et al. (2024c), and PerceptionTest Patraucean et al. (2023) include action-related questions, they primarily focus on general action recognition rather than fine-grained details. TemporalBench Cai et al. (2024) evaluates event-level action understanding using a different protocol that requires distinguishing correct captions from negative ones. Action-specific benchmarks, such as ActivityNet-Captions Krishna et al. (2017), TGIF-QA Jang et al. (2019), MoVid-Bench Chen et al. (2024b), EGOBody Zhang et al. (2022), and GRAB Taheri et al. (2020), address temporal sequences or human-object interactions but often overlook human-human interactions.
Our HAICBench provides a more comprehensive evaluation of human action understanding, covering five key aspects detailed in Table 1. In Appendix A, we compare HAICBench with three concurrent works MoVidBench Chen et al. (2024c), MotionBench Hong et al. (2025), and FAVORBench Tu et al. (2025).
@@ 2.3 Multi-modal Large Language Models
MLLMs Liu et al. (2023); Wang et al. (2024c, 2025a) generally process multi-modal input information and generate language output. Kosmos Huang et al. (2023) introduced an end-to-end framework that integrated visual inputs with LLM from cohesive training. Flamingo Alayrac et al. (2022) and InstructBLIP Dai et al. (2023) merged visual and linguistic features through cross-attention and a Q-Former module, respectively. MiniGPT-4 Zhu et al. (2024) and LLaVA Liu et al. (2023) simplified the integration by linearly projecting visual features directly into the LLM embedding space.
@@ 3.2 HAICTrain
We chose the WebVid-10M dataset Bain et al. (2021) as the video source for training due to its large scale and high vision quality. Initially, we apply the action video accumulation pipeline in Section 3.1 to collect action videos from the WebVid-10M. This process results in a collection of 126K videos, representing 1.2% of the original dataset. Then, we employ Gemini-1.5-Pro Team et al. (2024) to generate captions in the standardized format as depicted in Figure 1 referring to the videos and original captions. The specific prompts used for this process are detailed in Appendix B. We then employ an additional judgement to filter out failure cases that do not follow the pre-defined format or display low quality. This judgement ensures the quality of our data in HAICTrain.
@@ 3.3 HAICBench
Video Caption Annotation. To avoid potential overlap with public benchmarks, we choose the newly proposed Omega-multimodal dataset OMEGA Labs (2024) as our benchmark video source, which comprises over 30 million 2-minute video clips. We apply our video accumulation stage in Section 3.1 to obtain some video clips, and manually check these clips to obtain 1,140 video clips. Then, human annotators are required to describe the number of human subjects, noting static attributes (gender, age, clothing, accessories) for subject identification, followed by body movements, action sequences, and interactions with others in chronological order.
@@
QA Pair Generation. Based on captions above, we prompt GPT-4o Hurst et al. (2024) to generate multiple-choice question-answer pairs about human interactions, action details, action sequences, count, and human attributes. The prompts are presented in Appendix D. Each question-answer pair is checked by 2 annotators and will be corrected if there are any mistakes. All options are shuffled to avoid potential bias.
We get a total of 1,140 video-caption pairs and 7,548 QA pairs in total. Detailed statistics are presented in Figure 3. Although each video clip focuses on a single scene and is relatively short (less than 20 seconds), the captions are highly detailed, often exceeding 100 words. The word cloud analysis reveals that our captions provide comprehensive descriptions.
For evaluation, we construct the HAICBench-test (refered as HAICBench for short) by selecting 412 videos and 2,000 QA pairs, evenly distributed across all categories. Both HAICBench and the whole QA dataset will be released.
@@ 4 Experiments
Gemini-1.5-Pro Team et al. (2024) - 41.0 33.7 30.4 36.0 46.0 58.9
GPT-4o Hurst et al. (2024) 50* 40.0 37.8 30.6 35.8 44.6 51.4
VideoLLaMA2-7B Cheng et al. (2024) 64 21.0 15.7 18.9 17.3 28.9 24.2
LongVA-7B Zhang et al. (2024a) 64 19.3 14.7 13.8 11.5 36.0 20.3
InternVL2-8B Chen et al. (2024e) 64 28.8 31.0 19.9 21.5 31.5 40.2
Qwen2VL-7B Wang et al. (2024a) 64 27.0 18.5 22.1 19.9 35.7 38.9
LLaVA-Video-7B Zhang et al. (2024b) 64 29.2 23.1 15.8 23.7 35.0 48.3
LLaVA-Video-ActionPro-7B 64 35.2 27.5 16.9 30.7 44.2 56.9
@@
Gemini-1.5-Pro Team et al. (2024) - 67.0 71.3 52.1 73.5 56.7 81.4
GPT-4o Hurst et al. (2024) 50* 61.2 69.9 46.4 61.6 46.3 81.8
VideoLLaMA2-7B Cheng et al. (2024) 64 51.4 48.5 38.8 56.3 47.0 66.3
LongVA-7B Zhang et al. (2024a) 64 56.9 55.6 41.8 58.4 52.1 76.4
InternVL2-8B Chen et al. (2024e) 64 58.8 64.2 43.0 59.7 50.0 76.9
Qwen2VL-7B Wang et al. (2024a) 64 62.5 69.4 44.7 60.4 57.5 80.6
LLaVA-Video-7B Zhang et al. (2024b) 64 62.0 64.3 43.0 62.5 56.8 83.5
LLaVA-Video-ActionPro-7B 64 64.2 65.7 45.0 63.6 61.3 85.3
@@
LLaVA-Video 59.4 58.1 61.8 62.0
60.0 58.3 63.8 63.1
62.1 59.9 65.2 64.2
58.8 55.8 62.1 60.8
LLaVA-OneVision 57.8 47.9 57.1 59.4
59.3 55.7 61.8 60.7
60.6 56.9 62.7 61.3
58.8 55.8 60.4 59.8
@@ 4.1 Experimental Settings
Evaluation dataset. To evaluate human action understanding ability, we executed experiments on several action-related benchmarks, including MVBench, ActivityNet-QA, PerceptionTest, and our HAICBench. Note that we focused on questions related to human actions in these benchmarks. For MVBench, we performed a comparison of all sub-tasks whose names contain “Action”, resulting in 7 types: Action Antonym, Action Count, Action Sequence, Action Prediction, Action Localization, Fine-grained Action, and Unexpected Action. For PerceptionTest, we also selected all questions whose tags are related to action. More details are presented in Appendix C.
Evaluation metrics. For multiple-choice questions in MVBench, PerceptionTest, and HAICBench, we followed the prompting approach in LLaVA-1.5 Liu et al. (2024) and used accuracy as the metric. For ActivityNet-QA which features open-ended questions, we followed the evaluation protocol in Video-ChatGPT Maaz et al. (2024), utilizing GPT-4o-0513 to calculate accuracy. In terms of HAICBench, we implemented two evaluation settings: (1) Standard Evaluation, where the video and question are directly input into an MLLM to generate an answer; and (2) Caption Evaluation, where the model generates a caption for each video, and an LLM (Gemini-1.5-Pro) answers questions based on this caption, thereby assessing caption ability based on QA accuracy. To avoid the LLM guessing answers when captions do not mention related contents, we asked Gemini-1.5-Pro to refuse to answer those questions (see Appendix E).
Implementation Details. We used LLaVA-Video-7B as initial model and fine-tuned it on the training dataset. All parameters were updated during training. The model was fine-tuned for one epoch with a learning rate of 1e-5 for the LLM and 2e-6 for the vision encoder, with 256 batch size. We sampled 64 frames uniformly in both training and evaluation. During generation, we took the greedy search without randomness. Our experiments utilized 128 NVIDIA A800-80GB GPUs.